[figure: Bass; Tenor; Alt; Sopran; 4 4; 4 4; 4 4; 4 4; Je; Je; - su; - su,; mei; mei; -; ne; - ne; Freu; Freu; -; -; de,; - de,; mei; mei; -; nes; - nes; Her; Her; zens; zens; Wei; Wei; -; de,; - de,; 5; Je; Je; -; - su; su; mei; mei; -ne; - ne; Zier,; Zier,]

[figure: ach,; ach,; wie; wie; lang,; lang,; ach,; ach; lan; lan; -; -; ge; - ge; ist; ist; dem; dem; Her; Her; zen; -zen; 10; ban; ban; -; ge; - ge; und; und; ver; ver; - langt; - langt; nach; nach; dir!; dir!]

[figure: Got; Got; - tes; - tes; Lamm,; Lamm,; mein; mein; Bräu; Bräu; ti; ti; gam,; gam,; 15; au; au; ßer; -ßer; dir; dir; soll; soll; mir; mir; auf; auf; Er; Er; -; -; den; - den; nicht; nichts; sonst; sonst; lie; lie; bers; - bers; wer; wer; -; -; den.; - den.]